I need to see how to work the algebraic equation 3(x+7)-12=24
I do not just need the answer, but need to see how you got it. Thanks, Paula

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

3(x+7)-12=24
-------------
A better way to put it is, "How do I solve for x in the equation?", or "Solve the equation for x."
-----------------------
3(x+7)-12=24
Add 12 to both sides
3(x+7) = 36
Divide by 3
x+7 = 12
Subtract 7
x = 5
-----

Answer by richwmiller(17219)

I got it through sweat and tears.
Do you need to see the sweat and tears?